Please give a full explanation!

do a that is morted 3.5x 1o lnes per m Calcolak tie percentage uncertat in the numberイ【ines permetre ng @loche g↓ved fJA explanatimち
Answer #1

Generally, when the uncertainty is not quoted in an instrument, the error is assumed to carry 1 least significant digit.

in this case, n = 3.5 x 103

so, error will be, er = 0.1 x 103

therefore, n = 3.5\times 10^3 \pm 0.1\times 10^3 lines / m

now calculate the % error in this,

p = \frac{0.1\times 10^3}{3.5\times 10^3}\times 100 = 2.857%